abs(3-4x)+10=20
What is the answer and how you solve the answer?

Answer by edjones(8007)

abs(3-4x)+10=20
|3-4x|=10
.
3-4x=10
-4x=7
x=-7/4
.
3-4x=-10
-4x=-13
x=13/4
.
Ed